OpenStudy (anonymous):

Let f be defined by f(x)=−x^2−3x+5. f'(x)=-2x-3 An equation for the tangent line to y=f(x) at x=−1 is y=?

OpenStudy (anonymous):

i guess y comes out to be y=6-x

OpenStudy (anonymous):

you know the answer eq?

OpenStudy (anonymous):

how did u come out with y=6-x?

OpenStudy (anonymous):

at x=-1 y ll be y=7 right? like y= -(-1)^2-3(-1)+5=7

OpenStudy (anonymous):

and f'(x)=-1

OpenStudy (anonymous):

putting values in tangent eq (y-7)=(-1)(x-(-1))

OpenStudy (anonymous):

y=6-x.isn't it?

OpenStudy (anonymous):

that's right, but how did u know that t m=-1?

OpenStudy (anonymous):

o thats easy look slope has a point on function and they gave us a point x=-1 ok? put x=-1 in the eq f'(x) =-2x-3 when we put x=-1...f'(x)= -2(-1)-3=2-3=-1

OpenStudy (anonymous):

ooh i get it thank u

OpenStudy (anonymous):

welcome:)